A rectangular garden is 15 ft longer than it is wide. Its area is 1350 ft. What are its dimensions?

Answer by geetha_rama(94)

Let Length = y ft
Width = x ft
length is 15 ft longer than width
Hence y = x + 15
Area of the garden = xy
xy = 1350
x * (x + 15) = 1350
x^2 + 15x - 1350 = 0
x^2 + 45x - 30x -1350 =0
x(x+45) -30(x+45) = 0
(x+45)(x-30) = 0
x= -45, 30
Width has to be a positive number
Hence Width of the graden = 30 ft
Length of the garden = 45 ft